OpenStudy (anonymous):

can you please help me translate 48 divided by the sum of 3 and 9 into numerical expression

OpenStudy (unklerhaukus):

\[48 \div (3+9)\]

OpenStudy (anonymous):

\[48\div(3+9)\]

OpenStudy (anonymous):

ok thanks i have another if u could help with it also plz

OpenStudy (unklerhaukus):

sure

OpenStudy (anonymous):

and dt gives 4

OpenStudy (unklerhaukus):

\[+ \text{d}t =4\]

OpenStudy (anonymous):

translate the quotient of 12 plus 8 and 5 into numerical expression

OpenStudy (anonymous):

*that (48/12=4)

OpenStudy (unklerhaukus):

\[(12+8) \div 5\]

OpenStudy (anonymous):

how did u get that answer im confused

OpenStudy (anonymous):

then it also says to write result in both horizontel and vertical format????

OpenStudy (unklerhaukus):

well the quotient between two numbers ( a and b ) can be written as \[a\div b\]or equivalently\[\frac{a}{b}\]

OpenStudy (unklerhaukus):

so the first one can also be written as \[\frac{48} {3+9}\]

OpenStudy (unklerhaukus):

and the last one \[\frac{12+8}{5}\]

OpenStudy (anonymous):

thanks unkle rhaukus i finally get it

OpenStudy (unklerhaukus):

yay did you also know \[(a+b )\div c = \frac{a+b}{c} =\frac{a}{c} +\frac{b}{c}\]

OpenStudy (anonymous):

1 last question just check one i already did: the sum of twice four and nine

OpenStudy (unklerhaukus):

\[( 2\times 4) +9\]

OpenStudy (anonymous):

the last part yopu just did we havent quite got that far yet thanks i didnt do it right but i can see where i messed up

OpenStudy (anonymous):

it will be some what like this it's only a hint! Problem: Ms. Jensen likes to divide her class into groups of 2. Use mathematical symbols to represent all the students in her class. Solution: Let g represent the number of groups in Ms. Jensen's class. Then 2 · g, or 2g can represent "g groups of 2 students". In the problem above, the variable g represents the number of groups in Ms. Jensen's class. A variable is a symbol used to represent a number in an expression or an equation. The value of this number can vary (change). Let's look at an example in which we use a variable. Example 1: Write each phrase as a mathematical expression. Phrase Expression the sum of nine and eight 9 + 8 the sum of nine and a number x 9 + x The expression 9 + 8 represents a single number (17). This expression is a numerical expression, (also called an arithmetic expression). The expression 9 + x represents a value that can change. If x is 2, then the expression 9 + x has a value of 11. If x is 6, then the expression has a value of 15. So 9 + x is an algebraic expression. In the next few examples, we will be working solely with algebraic expressions. hope it helps!!